OpenStudy (anonymous):

If (x + 5)(x - 2) = 18, then which of the following statements is true? A. x + 7 = 0 or x - 4 = 0 B. x + 5 = 0 or x - 2 = 0 C. x + 5 = 6 or x - 2 = 3

OpenStudy (danjs):

distribute out the parenthesis first

OpenStudy (danjs):

x^2 -2x + 5x -10 = 18 x^2 + 3x - 28 = 0

OpenStudy (danjs):

now factor that into (x ) ( x ) = 0

OpenStudy (danjs):

7*4 = 28, so already can tell the first answer is it

OpenStudy (danjs):

(x +7)*(x - 4) = x^2 + 3x - 28 = 0 so (x+7) = 0 or (x-4) = 0

OpenStudy (anonymous):

Thank you for explaining. :)

OpenStudy (danjs):

welcome, anytime